OpenStudy (anonymous):

Is the ordered pair (1, -5) is a solution for the system x + 2y = -9 and x - y = 6. 1. Yes 2. No

OpenStudy (anonymous):

A yes

OpenStudy (anonymous):

Plug the value of x =1 in both equations. If both equations give back the value of y = -5 then (1, -5) is a solution

OpenStudy (anonymous):

For x + 2y = -9 a substitute for x =1 => 1 +2y = -9 => 2y = -10 => y = -5 and also for x - y = 6. substitute for x =1 => 1-y =6 => -y =5 => y = -5 Hence both equations satisfy ( 1, -5)

OpenStudy (anonymous):

Alternatively simplfy the system further to yield one equation: i.e. x + 2y = -9 (1) x - y = 6. (2) subtract equation (2) from equation (1) => x + 2y = -9 - (x - y = 6) => 3y = -15 => y = - 5 Now plug y = -5 in any one of the origal equations Thus if we choose equation (1) then x +2(-5) = -9 => x -10 = -9 => x =1 Hence the solution to the system of the equations is (1, -5)

OpenStudy (anonymous):

Just remember the ordered pai ( 1, -5) => when x = 1, the value of y = -5 Ordered pair is usually written as ( x, f(x)) = ( x, y)